# AutoOPE: Automated Off-Policy Estimator Selection

### v1

@@ Appendix B Off-policy Estimators Description
B.1 Model-Based Estimators
@@
The DM estimator [3] employs a supervised machine learning model $\hat{q}$ to estimate the mean reward $\mathbb{E}_{p(r_{t}|x_{t},a_{t})}\left[r_{t}\right]$. The policy value prediction is computed in this way:
@@
B.2 Model-Free Estimators
@@
The IPS estimator [29, 34] employs pure importance sampling to re-weight the average of rewards, accounting for the distribution shift between $\pi_{b}$ and $\pi_{e}$, since the average of rewards are collected under the logging policy, while the goal is to evaluate $\pi_{e}$. The policy value prediction is computed in this way:
@@
The SNIPS estimator [37, 14] tries to stabilize the IPS estimator normalizing the prediction by the empirical mean of the importance weights. The policy value prediction is computed in this way:
@@
The IPS-$\lambda$ estimator [26] shrink the importance ratio smoothly, with a sub-gaussian parameterized by a value $\lambda$ to reduce the IPS variance. The policy value prediction is computed in this way:
@@ B.3 Hybrid Estimators
The DR estimator [38, 12, 6] employs a reward model $\hat{q}$ like DM, but it improves the reward prediction by $\hat{q}$ using a correction term based on importance sampling. The policy value prediction is computed in this way:
@@
The SNDR estimator [38, 37, 14] is a variant of DR that shrinks the weights using the same normalization factor used in SNIPS. The policy value prediction is computed in this way:
@@
The DR-$\lambda$ estimator [26] is a variant of DR that shrinks the weights using the same smooth subgaussian by IPS-$\lambda$. The policy value prediction is computed in this way:
@@
The DRos estimator [35] is a variant of DR that reduces variance with an importance weight derived by minimizing the sharp bound of estimator MSE. The policy value prediction is computed in this way:
@@
The Switch estimator [44] is a variant of DR that switches between DM and DR, using DM when the importance ratio is below a threshold specified by the value of $\lambda$, and DR when $w(x_{t},a_{t})$ is above the threshold. The policy value prediction is computed in this way:
@@
Appendix C Meta-Dataset
C.1 Characteristics
As stated in Section 3, AutoOPE method is based on a zero-shot meta-learning approach framed as a supervised learning problem. In particular, the goal of this supervised learning problem is to predict, given an estimator $\hat{V}_{m}$ and an OPE task $\mathcal{T}_{\text{ope}}^{(j)}$ as input, the $MSE$ of $\hat{V}_{m}$ when applied to $\mathcal{T}_{\text{ope}}^{(j)}$.
Hence, to train a Random Forest regressor to predict this MSE, we need a finite set of possible estimators $\mathcal{V}$ such that $\hat{V}_{m}\in\mathcal{V}$ defined a-priori, and a large set of OPE tasks $\mathcal{M}$. The set $\mathcal{M}$ is a collection of synthetically generated OPE tasks, obtained from a data generator function $G$, which we describe in further detail in Section C.2.
The training dataset for the Random Forest is obtained extracting suitable input features from the estimator $\hat{V}_{m}$ and an OPE task $\mathcal{T}_{\text{ope}}^{(j)}$, that the Random Forest can work with. This is done using a feature extraction function $g$. Then, the target is the ground-truth $MSE$ of $\hat{V}_{m}$ when applied to $\mathcal{T}_{\text{ope}}^{(j)}$. The training dataset can be expressed as: $\{g(\mathcal{V}\times\mathcal{M}),\ MSE(\mathcal{V}\times\mathcal{M})\}$
The meta-learning dataset $\mathcal{M}$ contains $250,000$ OPE tasks, each associated to all the candidate estimators in $\mathcal{V}$ that are $21$, for a total of $5.2$ million samples. Note that we removed the few trivial OPE tasks which had all rewards equal to 0 or 1.
@@ C.1 Characteristics
The set $\mathcal{V}$ of candidate estimators includes several model-free, model-based and hybrid estimators, as done in [40]. Estimators hyperparameters, when present, were selected using the SLOPE method [36]. All the estimators used are listed in Table 3, and a brief explanation of each of them is reported in appendix B.
Each of the six estimators with a reward model is employed in conjunction with three reward predictors: Random Forest Classifier and Logistic Regression from the Scikit-Learn [^4] Scikit-Learn available with 3-Clause BSD license: https://github.com/scikit-learn/scikit-learn library [28], and LGBM Classifier from the Python package of LightGBM [^5] LightGBM available with MIT license: https://github.com/microsoft/LightGBM/tree/master [16]. The reward models were fitted using a 3-fold cross-fitting procedure to reduce possible overfitting [27].
@@
Table 3: List of the OPE estimators used in this paper. See Appendix B for their description.
Features
We compute 43 features for each OPE task that can be categorized into three groups:
  • Policies-independent features, that describe structural properties of the CB scenario and are unrelated from the specific $\pi_{b},\ \pi_{e}$, e.g., the number of rounds, the reward mean and variance, the number of actions.
  • Policies-dependent features, such as the dissimilarity between the two policies $\pi_{b},\ \pi_{e}$, measured with different geometrical and statistical distances.
  • OPE estimator features, that refer to specific properties of the OPE estimator used in that task, e.g., if the estimator is self-normalized, uses importance sampling, employs a reward model, and if so, which one.
The comprehensive list of features, together with their mathematical formulations, is reported in Appendix E.
@@
The target variable is the MSE between the prediction of an estimator $\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(j)})\ \text{with}\ \hat{V}_{m}\in\mathcal{V}$ and the on-policy estimated ground-truth policy value $V(\pi_{e})$. Indeed, together with each synthetic OPE task $\mathcal{T}_{\text{ope}}^{(j)}$, we also generate a ground-truth dataset, denoted as $\mathcal{D}_{gt}^{(j)}\coloneqq\left\{(x_{t},a_{t},r_{t},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}})\right\}_{t}^{n_{gt}}$, that is a large CB dataset collected under the evaluation policy $\pi_{e}$.
Realizations
In order to improve the robustness of the data, both the features and MSE are computed as the average of $n_{gen}$ realizations $\mathcal{T}_{ope}^{(j)}(s)$ of each $\mathcal{T}_{ope}^{(j)}$. The features and the target MSE are computed for each candidate estimator and averaged across all realizations to obtain the final values used in the meta-learning dataset.
C.2 Generation
@@
We can formalize the generation of the pair $(\mathcal{T}_{\text{ope}}^{(j)},\mathcal{D}_{gt}^{(j)})$ as the output of a function $G$ which serves as synthetic data generator, so $(\mathcal{T}_{\text{ope}}^{(j)},\mathcal{D}_{gt}^{(j)})=G(p_{1}^{(j)},\dots,p_{P}^{(j)})$, with $(p_{1}^{(j)},\dots,p_{P}^{(j)})$ drawn from $\Phi=(\phi_{1}(\cdot),\dots,\phi_{P}(\cdot))$. For each new OPE task, we sample a value for each parameter $p_{i}$ from the respective distribution $\phi_{i}$, and generate the datasets pair. Function $G$ is highly configurable since each parameter has a distribution that can be specified, which allows to represent a wide range of scenarios. Among the parameters $\left\{p_{i}\right\}_{i}^{P}$ that change the OPE task characteristics we include: the evaluation and logging policies, the number of possible actions, the number of rounds in the logging data, the context dimension, the reward distribution.
Below we provide the complete list of these parameters, along with their distributions $\left\{\phi_{i}\right\}_{i}^{P}$, providing insight into the characteristics considered for OPE tasks:
@@
  • $n_{data}$: a positive integer determining the number of synthetic OPE task data ($\mathcal{T}_{\text{ope}}$) to generate. In total, we generated $n_{data}=250,000$ different OPE task.
@@
Introduced the parameters used in the generation process, the procedure for obtaining the two datasets $(\mathcal{T}_{\text{ope}},\mathcal{D}_{gt})$ can be formally described.
The procedure is outlined by Algorithm 3. Internally to the function Generate, called in the Algorithm 3:
@@
Afterwards, within the function GenerateOPE, called in the Algorithm 3:
@@
These data are sufficient to formally define $\mathcal{T}_{\text{ope}}$ as in Equation 1.
The procedure to generate $\mathcal{T}_{\text{ope}}$ is repeated $n_{gen}$ times, obtaining $n_{gen}$ realizations of the $\mathcal{T}_{\text{ope}}$ dataset.
@@
0:  number of OPE task generations $n_{data}$, number of realizations $n_{gen}$ of each OPE dataset, a big number of rounds $n_{gt}$ to sample under $\pi_{e}$ for the evaluation dataset, space of generation parameters $\Phi$
0:  Collection of OPE datasets and evaluation datasets $\left\{\{\mathcal{T}^{(j)}_{ope}(s)\}_{s}^{n_{gen}},\ \mathcal{D}^{(j)}_{gt}\right\}_{j=1}^{n_{data}}$
1:  for $j=1,\ldots,n_{data}$ do
2:     $(n_{a},n_{b},d_{x},q,\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}})\sim\Phi$
3:     $\mathcal{A}\leftarrow\left\{0,\dots,n_{a}-1\right\}$
4:     $\mathcal{D}^{(j)}_{gt}\leftarrow\mathrm{Generate}(\mathcal{A},d_{x},\beta_{e},f_{\pi_{e}},q,n_{gt})$
5:     for $s=1,\ldots,n_{gen}$ do
6:        $\mathcal{T}^{(j)}_{ope}(s)\leftarrow\mathrm{GenerateOPE}(\mathcal{A},d_{x},\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}},q,n_{b},s)$
7:     end for
8:  end for
Appendix D Supervised Model Training
@@
As a supervised model, we select the Random Forest regressor provided by the scikit-learn package [^7] https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.RandomForestRegressor.html.
@@
In order to train and evaluate the AutoOPE supervised model, we split the meta-learning dataset in 80% for optimization and training, and 20% for testing, so that the testing data never contains OPE tasks the supervised model was trained on.
@@
AutoOPE uses a Random Forest as a supervised model. In order to select the best hyperparameters we rely on a Bayesian Search optimizing the 5-fold cross-validation Regret between the predicted MSE of the OPE estimator selected by AutoOPE, $\hat{V}_{\hat{m}}$, and the best estimator according to the MSE on the ground-truth dataset. We define the Regret as:
$\text{Regret}\coloneqq MSE(\hat{V}_{\hat{m}}(\pi_{e}))-MSE(\hat{V}_{m^{*}}(\pi_{e}))\ ,$ (4)
This means that at each step of the Bayesian Search, the training dataset is split into five folds, and a Random Forest regression model (with fixed hyperparameters) is fitted for each of the possible five subsets of four data folds, resulting in five models in total. Each model is then evaluated on the remaining out-of-bag fold in terms of Regret.
The Regret values are averaged over the different five models, and the 5-fold cross-validation metric is obtained, that is the function optimized by the Bayesian Search.
The Bayesian Search is implemented on top of the Scikit-Optimize [^8] Scikit-Optimize available with 3-Clause BSD license: https://github.com/scikit-optimize/scikit-optimize/tree/master library, and involves Bayesian optimization using Gaussian Processes to approximate the objective function, making the assumption that function points are distributed as a multivariate Gaussian. The Bayesian Optimization is executed on the training data, with 100 calls made to optimize the hyperparameters.
@@
Appendix E Features Design
This Appendix mathematically describes the input features for the Random Forest in AutoOPE. They are:
@@ Appendix E Features Design
    • Neyman chi-squared distance between $\pi_{b},\pi_{e}$, which we included because it is equivalent (up to a constant) to the exponentiated Rényi divergence (shown to be an important data characteristic by Metelli et al. [24]):
@@
In Figure 4 is reported the complete Feature Importance plot, measured with Mean Decrease in Impurity (MDI).
@@
Appendix F Additional Experiments and Ablation Studies
F.1 Scaling Experiment

### v2

@@ Appendix C OFF-POLICY ESTIMATORS DESCRIPTION
C.1 Model-Based Estimators
@@
The DM estimator (Beygelzimer and Langford, 2009b) employs a supervised machine learning model $\hat{q}$ to estimate the mean reward $\mathbb{E}_{p(r_{t}|x_{t},a_{t})}\left[r_{t}\right]$. The policy value prediction is computed in this way:
@@
C.2 Model-Free Estimators
@@
The IPS estimator (Precup et al., 2000; Strehl et al., 2010) employs pure importance sampling to re-weight the average of rewards, accounting for the distribution shift between $\pi_{b}$ and $\pi_{e}$, since the average of rewards are collected under the logging policy, while the goal is to evaluate $\pi_{e}$. The policy value prediction is computed in this way:
@@
The SNIPS estimator (Swaminathan and Joachims, 2015; Kallus and Uehara, 2019) tries to stabilize the IPS estimator normalizing the prediction by the empirical mean of the importance weights. The policy value prediction is computed in this way:
@@
The IPS-$\lambda$ estimator (Metelli et al., 2021) shrink the importance ratio smoothly, with a sub-gaussian parameterized by a value $\lambda$ to reduce the IPS variance. The policy value prediction is computed in this way:
@@ C.3 Hybrid Estimators
The DR estimator (Thomas and Brunskill, 2016; Jiang and Li, 2016; Dudík et al., 2015) employs a reward model $\hat{q}$ like DM, but it improves the reward prediction by $\hat{q}$ using a correction term based on importance sampling. The policy value prediction is computed in this way:
@@
The SNDR estimator (Thomas and Brunskill, 2016; Swaminathan and Joachims, 2015; Kallus and Uehara, 2019) is a variant of DR that shrinks the weights using the same normalization factor used in SNIPS. The policy value prediction is computed in this way:
@@
The DR-$\lambda$ estimator (Metelli et al., 2021) is a variant of DR that shrinks the weights using the same smooth subgaussian by IPS-$\lambda$. The policy value prediction is computed in this way:
@@
The DRos estimator (Su et al., 2020a) is a variant of DR that reduces variance with an importance weight derived by minimizing the sharp bound of estimator MSE. The policy value prediction is computed in this way:
@@
The Switch estimator (Wang et al., 2017) is a variant of DR that switches between DM and DR, using DM when the importance ratio is below a threshold specified by the value of $\lambda$, and DR when $w(x_{t},a_{t})$ is above the threshold. The policy value prediction is computed in this way:
@@
Appendix D SYNTHETIC DATASET
D.1 Characteristics
As stated in Section 3, AutoOPE method is based on a zero-shot approach framed as a supervised learning problem. In particular, the goal of this supervised learning problem is to predict, given an estimator $\hat{V}_{m}$ and an OPE task $\mathcal{T}_{\text{ope}}^{(j)}$ as input, the MSE of $\hat{V}_{m}$ when applied to $\mathcal{T}_{\text{ope}}^{(j)}$.
Hence, to train a Random Forest regressor to predict this MSE, we need a finite set of possible estimators $\mathcal{V}$ such that $\hat{V}_{m}\in\mathcal{V}$ defined a-priori, and a large set of OPE tasks $\mathcal{M}$. The set $\mathcal{M}$ is a collection of synthetically generated OPE tasks, obtained from a data generator function $G$, which we describe in further detail in Section D.2.
The training dataset for the Random Forest is obtained extracting suitable input features from the estimator $\hat{V}_{m}$ and an OPE task $\mathcal{T}_{\text{ope}}^{(j)}$, that the Random Forest can work with. This is done using a feature extraction function $g$ (described in Appendix F.). Then, the target is the ground-truth MSE of $\hat{V}_{m}$ when applied to $\mathcal{T}_{\text{ope}}^{(j)}$. The training dataset can be expressed as: $\{\mathrm{input:}\ g(\mathcal{V}\times\mathcal{M}),\ \mathrm{target:}\ \text{MSE}(\mathcal{V}\times\mathcal{M})\}$
The meta-dataset $\mathcal{M}$ contains $250,000$ OPE tasks, and for each task $\mathcal{T}_{\text{ope}}^{(j)}$, $n_{gen}=10$ realizations $\{\mathcal{T}_{\text{ope}}^{(j)}(s)\}_{s}^{n_{gen}}$ are produced, for a total of $2,500,000$ OPE tasks. Each task is associated to all the candidate estimators in $\mathcal{V}$ that are $21$, for a total of $52.5$ million samples. Note that we removed the few trivial OPE tasks which had all rewards equal to 0 or 1.
@@ D.1 Characteristics
The set $\mathcal{V}$ of candidate estimators includes several model-free, model-based and hybrid estimators, as done in (Udagawa et al., 2023). Estimators hyperparameters, when present, were selected using the SLOPE method (Su et al., 2020b). All the estimators used are listed in Table 3, and a brief explanation of each of them is reported in appendix C.
Each of the six estimators with a reward model is employed in conjunction with three reward predictors: Random Forest Classifier and Logistic Regression from the Scikit-Learn [^5] Scikit-Learn available with 3-Clause BSD license: https://github.com/scikit-learn/scikit-learn library (Pedregosa et al., 2011), and LGBM Classifier from the Python package of LightGBM [^6] LightGBM available with MIT license: https://github.com/microsoft/LightGBM/tree/master (Ke et al., 2017). The reward models were fitted using a 3-fold cross-fitting procedure to reduce possible overfitting (Narita et al., 2021).
@@
Table 3: List of the OPE estimators used in this paper. See Appendix C for their description.
@@
The target variable is the MSE between the prediction of an estimator $\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(j)}(s))\ \text{with}\ \hat{V}_{m}\in\mathcal{V}$ and the on-policy estimated ground-truth policy value $V(\pi_{e})$. Indeed, together with each synthetic OPE task $\mathcal{T}_{\text{ope}}^{(j)}(s)$, we also generate a ground-truth dataset, denoted as $\mathcal{D}_{gt}^{(j)}\coloneqq\left\{(x_{t},a_{t},r_{t},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}})\right\}_{t}^{n_{gt}}$, that is a large CB dataset collected under the evaluation policy $\pi_{e}$, which is used to compute the ground-truth MSE, by averaging the collected rewards (see Appendix J for further details).
D.2 Generation
@@
Synthetizing the previous section, we can say that it is needed a pair $(\mathcal{T}_{\text{ope}}^{(j)}(s),\mathcal{D}_{gt}^{(j)})$ and one estimator $\hat{V}_{m}$, to generate one row of the dataset on which the Random Forest will be trained. We can formalize the generation of the pair ($(\{\mathcal{T}_{\text{ope}}^{(j)}(s)\}_{s}^{n_{gen}},\mathcal{D}_{gt}^{(j)})$ as the output of a function $G$ which serves as synthetic data generator, so $(\mathcal{T}_{\text{ope}}^{(j)}(s),\mathcal{D}_{gt}^{(j)})=G(p_{1}^{(j)},\dots,p_{P}^{(j)})$, with the parameters $(p_{1}^{(j)},\dots,p_{P}^{(j)})$ drawn from $\Phi=(\phi_{1}(\cdot),\dots,\phi_{P}(\cdot))$. For each new OPE task, we sample a value for each parameter $p_{i}$ from the respective distribution $\phi_{i}$, and generate the datasets pair. Function $G$ is highly configurable since each parameter has a distribution that can be specified, which allows to represent a wide range of scenarios. Among the parameters $\left\{p_{i}\right\}_{i}^{P}$ that change the OPE task characteristics we include: the evaluation and logging policies, the number of possible actions, the number of rounds in the logging data, the context dimension, the reward distribution.
Below we provide the complete list of these parameters, along with their distributions $\left\{\phi_{i}\right\}_{i}^{P}$, providing insights into the characteristics considered for OPE tasks:
@@
Introduced the parameters used in the generation process, the procedure for obtaining the pair ($(\{\mathcal{T}_{\text{ope}}^{(j)}(s)\}_{s}^{n_{gen}},\mathcal{D}_{gt}^{(j)})$ can be formally described.
The procedure is outlined by Algorithm 3, which internally calls the function Generate and the function GenerateOPE, explained in the following:
Generate
@@
  5. (5)
    The data produced are sufficient to formally define a logging dataset to compute the on-policy value estimate as rewards average.
GenerateOPE
@@
  5. (5)
    The data produced are sufficient to formally define $\mathcal{T}_{\text{ope}}^{(j)}(s)$ as in Equation 1.
@@
0:  number of realizations $n_{gen}$ of each OPE dataset, a big number of rounds $n_{gt}$ to sample under $\pi_{e}$ for the evaluation dataset, space of generation parameters $\Phi$
0:  Collection of OPE datasets and evaluation datasets $\{\mathcal{T}^{(j)}_{ope}(s)\}_{s}^{n_{gen}},\ \mathcal{D}^{(j)}_{gt}\}_{j=1}^{n_{data}}$
1:  $(n_{a},n_{b},d_{x},q,\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}})\sim\Phi$
2:  $\mathcal{A}\leftarrow\left\{0,\dots,n_{a}-1\right\}$
3:  $\mathcal{D}^{(j)}_{gt}\leftarrow\mathrm{Generate}(\mathcal{A},d_{x},\beta_{e},f_{\pi_{e}},q,n_{gt})$
4:  for $s=1,\ldots,n_{gen}$ do
5:     $\mathcal{T}^{(j)}_{ope}(s)\leftarrow\mathrm{GenerateOPE}(\mathcal{A},d_{x},\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}},q,n_{b},s)$
6:  end for
Appendix E SUPERVISED MODEL TRAINING
@@
As a supervised model, we select the Random Forest regressor provided by the scikit-learn package [^8] https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.RandomForestRegressor.html. The learning pipeline comprehend also some preprocessing steps and transformations:
  • Categorical features are one-hot encoded.
  • Numerical features are clipped to a maximum value of $10^{10}$.
  • The function $log(1+x)$ is applied to positive features with a skewness higher than 1 (and to the target variable).
  • Each numerical feature is scaled to the interval $(0,1)$ if the feature is positive and to $(-1,1)$ if can assume negative values. The scaling is applied also the target variable.
@@
In order to train and evaluate the AutoOPE supervised model, we split the meta-dataset into three parts, 60% for training, 20% for validation, and 20% for testing.
@@
AutoOPE uses a Random Forest as a supervised model. In order to select the best hyperparameters we rely on a Bayesian Search optimizing the Regret between the ground-truth MSE of the OPE estimator selected by AutoOPE, $\hat{V}_{\hat{m}}$, and the best estimator according to the MSE on the ground-truth dataset. We define the Regret as:
$\text{Regret}\coloneqq\text{MSE}(\hat{V}_{\hat{m}}(\pi_{e}))-\text{MSE}(\hat{V}_{m^{*}}(\pi_{e}))\ ,$ (5)
This means that at each step of the Bayesian Search, a Random Forest regression model (with fixed hyperparameters) is fitted on the training set, and then is evaluated on the validation set in terms of Regret. This Regret is the objective function that the Bayesian Search tries to optimize. Finally, the Random Forest is trained on the union of the training and validation set with the best hyperparameters found by the Bayesian Search, and is evaluated on the test set to asses the performance.
The Bayesian Search is implemented on top of the Scikit-Optimize [^9] Scikit-Optimize available with 3-Clause BSD license: https://github.com/scikit-optimize/scikit-optimize/tree/master library, and involves Bayesian optimization using Gaussian Processes to approximate the objective function, making the assumption that function points are distributed as a multivariate Gaussian. For our model, 50 Bayesian Search iterations are performed.
@@
Appendix F FEATURE DESIGN
This Appendix mathematically describes the input features computed by the feature extractor $g$. They are:
@@ Appendix F FEATURE DESIGN
    • Neyman chi-squared distance between $\pi_{b},\pi_{e}$, which we included because it is equivalent (up to a constant) to the exponentiated Rényi divergence (shown to be an important data characteristic by Metelli et al. (2018)):
@@
Feature Importance
In Figure 4 is reported the complete Feature Importance plot, measured with Mean Decrease in Impurity (MDI). The 10 policy-independent features appear to be of limited importance and correspond to 5 of the 15 least important features. Only 2 policy-independent features belong to the first 15 most important features. Policy-dependent features, which are 24, are generally useful for the model, with 9 among the 15 most important features and 6 out of 24 among the 15 least important. The KL divergence is the most important feature, which confirms what previous studies in theory have found (Chatterjee and Diaconis, 2018). The 9 OPE estimator features are similarly spread, with 4 among the most important features and 4 among the least important ones.
The insights gained from this feature importance analysis is another contribution. In particular, it can shed light on future theoretical findings, lead by analyzing the features that had a significant impact on the model’s performance.
@@
Appendix G ADDITIONAL EXPERIMENTS
G.1 Additional Experiments on CIFAR-10
The CIFAR-10 dataset (Krizhevsky et al., 2009) is a widely-used benchmark for image classification tasks, with mutually exclusive labeling.
CB Data
To adapt the CIFAR-10 dataset for CB we perform the standard supervised to bandit conversion (Beygelzimer and Langford, 2009a; Joachims et al., 2018), reinterpreting the input features (i.e., images pixels scaled in $(0,1)$) as context features and the class labels as actions. Furthermore, we partition the dataset into two subsets, the first one is used as logging data $\mathcal{D}_{b}$ for the OPE task, while the second is used to generate both logging and evaluation policies. In order to generate the policies, the second data split is used to train two Logistic Regression classifiers. Each classifier defines a deterministic policy $\pi_{det}$ because, given a context, it predicts a specific class label (action) with a probability of 1, while all other labels having a probability of 0. To produce the final stochastic policies we blend the deterministic policy with a uniform random policy $\pi_{u}(a|x)$ using using parameters $\alpha_{b}$, $\alpha_{e}$:
[…]
Results
G.2 Additional Synthetic Experiments
In this section we describe all the experiments on synthetic data, following the experimental pipeline of (Udagawa et al., 2023).
CB Data
The synthetic logging data is generated using the following procedure and distributions (same procedure used in (Udagawa et al., 2023)):
  • A total of 10 possible actions is considered.
As already said, we divide the synthetic experiments into two main settings: one named Logging 1, where we set $\beta_{1}=2$ and $\beta_{2}=-2$, and the other called Logging 2, where $\beta_{1}=3$ and $\beta_{2}=7$, in line with the experimental analysis performed by Udagawa et al. (2023).
To provide precise performance assessments, in all the 21 experiments of both Logging 1 and 2 we generate $n_{data}=100$ realizations of the logging dataset, as done in the procedure outlined in (Udagawa et al., 2023). The performance metrics are subsequently averaged over all realizations to yield more robust performance estimates with their confidence intervals.
Results
The results obtained from Logging 1 are reported in Figure 7 while those of Logging 2 in Figure 7. Since these synthetic experiments are the same of (Udagawa et al., 2023), we use their optimal values for the other hyperparameters such as the network structure, regularization, learning rate. In both cases our proposed AutoOPE outperforms PAS-IF across almost all evaluation policies in terms of Spearman’s rank correlation coefficient. In terms of Relative Regret, AutoOPE performs similar to PAS-IF in the first half settings of the Logging 1 experiments. However, for high values of $\beta_{e}$, PAS-IF slightly outperform AutoOPE. In the Logging 2 set of experiments, we see that for all the $\beta_{e}\geq 0$ AutoOPE outperforms PAS-IF, but PAS-IF performs better in the settings with $\beta_{e}<0$. The fact that AutoOPE exhibits better Spearman’s Rank Correlation indicates that, while both methods predict similar values for the best estimator, AutoOPE is more consistent in ranking all of them.
Appendix H ABLATION STUDIES
In this appendix, we report some ablation studies. Due to the specific settings of these experiments, several new model optimizations need to be executed. Given the limited time and resources, we consider a smaller version of the meta-dataset $\mathcal{M}$ to produce results within a reasonable timeframe. For the same reason, we fix the number of trees in the Random Forest to 200, which we find to be a reasonable trade-off between power and manageability.
In the original dataset, each OPE task generated has 10 realizations, with a row of feature values for each realization. In this smaller version of the meta-dataset, however, we consider only one row per OPE task, obtained by averaging the features of the 10 realizations.
These experiments aim to demonstrate that limiting certain characteristics of the meta-dataset’s OPE tasks, such as the number of actions or the KL-divergence of the policies, can affect model performance. Importantly, the conclusions obtained from the following experiments remain valid, independently of the meta-dataset size.
H.1 Scaling Experiment